The sum of the first 3 terms of geometric sequence is 189 and the sum of the first 6 terms is 182.


1
Expert's answer
2022-02-14T16:56:06-0500
"b_1+b_1r+b_1r^2=189"

"b_1+b_1r+b_1r^2+b_1r^3+b_1r^4+b_1r^5=182"



"b_1(1+r+r^2)=189"

"b_1r^3(1+r+r^2)=-7"

"r^3=-\dfrac{7}{189}"

"r=-\dfrac{1}{3}"

"b_1(1+(-\dfrac{1}{3})+(-\dfrac{1}{3})^2)=189"

"b_1=\dfrac{189(9)}{7}"

"b_1=243"

"b_1=243, r=-\dfrac{1}{3}"